### v2

# DreamRunner: Fine-Grained Storytelling Video Generation with Retrieval-Augmented Motion Adaptation

DreamRunner: Fine-Grained Storytelling Video Generation with Retrieval-Augmented Motion Adaptation
Zun Wang   Jialu Li   Han Lin   Jaehong Yoon   Mohit Bansal
UNC Chapel Hill
{zunwang,jialuli,hanlincs,jhyoon,mbansal}@cs.unc.edu
https://dreamrunner-story2video.github.io/
Appendix A Ablations of RAG pipeline
We provide additional experiments in Section 3.2 to demonstrate the effectiveness of our data processing approach within the retrieval pipeline. Table 1 primarily evaluates the impact of the maximal number of retrieved videos and the use of CLIP and ViCLIP for filtering. For efficiency, we evaluate a subset of eight motions selected from the full pool of 64 motions. The results indicate that, compared to the CogVideoX zero-shot baseline (Row 1), retrieving videos without any filtering (Row 2) improves performance, even though BM-25 [robertson2009probabilistic] retrieval introduces some noise. This highlights the importance of retrieval itself. Furthermore, adding CLIP and ViCLIP as filters (Row 4) further enhances performance, showcasing the benefit of using semantically aligned videos for improved motion learning. Additionally, retrieving a sufficient number of videos is critical, as evidenced by Row 3, where limiting retrieval to a maximum of three videos results in poorer performance compared to retrieving 20 videos (Row 4).
  Max. #Retrieval CLIP+ViCLIP filter CLIP ViCLIP
0 $\times$× 23.42 20.56
20 $\times$× 24.01 22.51
3 24.45 22.80
20 25.47 23.66
Table 1: Pipeline component ablation on retrieval-augmented test-time adaptation for learning a better motion prior.
Appendix B Evaluation Metrics
We provide detailed evaluation metrics for evaluating generated storytelling videos in multiple dimensions in Section 4.1.
  • Similarity to Reference Images: We assess the alignment between the generated videos and reference images using Image-Image CLIP [hessel2021clipscore] and DINO [oquab2023dinov2] scores. The evaluation is conducted by averaging the CLIP/DINO similarity scores with each frame of the generated video and the reference images, with CLIP-L14 (336px) [hessel2021clipscore] and DINOv2-B16 [oquab2023dinov2] as image encoders.
  • Similarity to Full Narration per scene: We measure the alignment between the full narration and the generated videos for each scene using Image-Text CLIP [hessel2021clipscore] and Video-Text ViCLIP [wang2023internvid] scores. For the CLIP score, we uniformly sample eight frames from the single-scene video and compute the average score between each frame and the full narration. For the ViCLIP score, we directly use the alignment score between the video and the narration. Per-scene scores are averaged to obtain the overall score for the multi-scene storytelling video.
  • Fine-Grained Text Alignment per Scene: We also assess fine-grained alignment to textual descriptions using Image-Text CLIP [hessel2021clipscore] and Video-Text ViCLIP [wang2023internvid] scores. In our story, each narration contains two motions, which we decouple into two consecutive single-motion descriptions using an LLM [openai_gpt4o]. For each generated single-scene video, we divide it into two segments at the temporal midpoint. We then compute the CLIP/ViCLIP scores between the first segment and the first description and between the second segment and the second description. The average of these two scores provides the per-scene score, and the per-scene scores are further averaged to compute the overall score for the storytelling video.
  • Transition: We assess whether the single-scene video achieves smooth transitions between two motions. To evaluate this, we uniformly sample four frames from the video per scene and calculate the average DINO similarity between adjacent frames. A higher transition score indicates smoother transitions, as it reflects minimal changes in the background across frames.
Appendix C Compositional T2V Examples
In this section, we present qualitative examples demonstrating the capabilities of DreamRunner in compositional text-to-video generation. DreamRunner effectively generates videos with accurate action binding to different characters, consistent attributes across objects, dynamic attribute changes, motion control, object interactions, and appropriate spatial relationships between objects.
For action binding, as shown in Figure 1, DreamRunner generates distinct motions for two objects (e.g., a wolf howling into a microphone and a fox playing the drums), ensuring the actions are correctly bound to their respective objects without interference.
For consistent attribute binding, as illustrated in Figure 2, our approach maintains separate attributes for different objects (e.g., a spherical globe and a cube-shaped clock) without any overlap or inconsistency.
For dynamic attribute changes, as shown in Figure 3, DreamRunner naturally transitions object attributes over time (e.g., metal gradually rusting throughout the video).
For motion control, as depicted in Figure 4, DreamRunner successfully directs object movements in different trajectories (e.g., a kite moving from left to right and a cyclist traveling from right to left).
For object interactions, as illustrated in Figure 5, interactions are accurately modeled, adhering to physical rules (e.g., pottery shattering upon hitting the floor).
Finally, for spatial relationships, as shown in Figure 6, DreamRunner generates rare or imaginative configurations (e.g., a duck positioned under a spacecraft) while maintaining spatial coherence.
These examples highlight the strong performance of DreamRunner in generating high-quality compositional text-to-video outputs.
Appendix D Characters.
We provide character examples in Figure 7, where the first four are generated using FLUX [flux], and the others are collected from existing customization datasets (CustomConcept101 [kumari2023multi] and Dreambooth [ruiz2023dreambooth]).
Appendix E Single-Character Examples
In this section, we present qualitative examples of video generation featuring a single main character. As shown in Figure 9 and Figure 8, DreamRunner generates consistent characters throughout the entire story. Additionally, in each scene, DreamRunner effectively captures multiple events, such as the mermaid first wandering through the plants and then examining unique shells.
Appendix F Multi-Character Examples
In this section, we present qualitative examples of video generation featuring multiple characters. As illustrated in Figure 10 and Figure 11, DreamRunner generates multi-scene, multi-character videos where each character retains its own motion and interacts seamlessly with others, without any interference. For instance, in Figure 10, the witch is shown pouring ingredients while the cat wanders around the room. Even as the cat approaches the witch, their motions remain independent, and the appearance of both characters is consistently preserved throughout the scene.
Appendix G Region-Based 3D Attention Masks
Figure 12 shows and example of the attention Mask of our Spatial-Temporal Region-based 3D Attention introduced in Section LABEL:sec:subsec:attn. Different text colors represent different conditions, while the white region indicates the masked areas. Note that for simplicity, we reduce each condition to two words, each frame to three segments, and display only three conditions and two frames in the figure. In practice, conditions can be longer and more numerous, frames can have more segments, and there are 12 latent frames in total.
Appendix H LLM Prompts
We provide detailed LLM prompts for both high-level plans and fine-grained plans in Listing LABEL:lst:hlplan and Listing LABEL:lst:fgplan. respectively. For high-level plans, we use a simple in-context example with instructions, while fine-grained plans require reasoning before generating the output. Example outputs are shown in Listings LABEL:lst:hleg and LABEL:lst:fgeg. For multi-character scenarios, we modify character-related words and examples and limit required motions to a maximum of four.
Figure 1: Qualitative results of DreamRunner generated with prompts characterizing action binding.
Figure 2: Qualitative results of DreamRunner generated with prompts characterizing consistent attribute binding.
Figure 3: Qualitative results of DreamRunner generated with prompts characterizing dynamic attribute binding.
Figure 4: Qualitative results of DreamRunner generated with prompts characterizing motion binding.
Figure 5: Qualitative results of DreamRunner generated with prompts characterizing object interactions.
Figure 6: Qualitative results of DreamRunner generated with prompts characterizing spatial relationships.
Figure 7: Qualitative results of DreamRunner generated with a single character (mermaid).
Figure 8: Qualitative results of DreamRunner generated with a single character (astronaut).
Figure 9: Qualitative results of DreamRunner generated with a single character (mermaid).
Figure 10: Qualitative results of DreamRunner generated with multiple characters (witch and cat 1).
Figure 11: Qualitative results of DreamRunner generated with multiple characters (warrior and dog 2).
Figure 12: Visualization of spatial-temporal region-based 3D attention masks. Different text colors represent different conditions, while the white region indicates the masked areas. For simplicity, we reduce each condition to two words, each frame to three segments, and display only three conditions and two frames in the figure. In practice, conditions can be longer and more numerous, frames can have more segments, and there are 12 latent frames in total.
Listing 1: The LLM prompt for high-level planning (Section 3.1, Story-Level Croase-Grained Planning). User may input the high-level planning to generate at the [Input] highlighted in blue. We highlight scene, motions, narrations using same colors as Section 3.1.
Consider you are an expert in writing stories. I will provide you with a topic, and you need to create a multi-scene story with 5 to 8 scenes. Each scene should describe the events taking place, emphasizing highlighting human actions or motions. Limit each scene to a maximum of 2 distinct human motions. Your output should include both the scene and motions as well as the narration, where the motion should be in present progressive.
[Input] Teddys one day
[Output]
Scene 1: bedroom
Motions:
waking up, stretching arms
Narration:
Teddy starts his day in his bedroom. He wakes up from the bed and stretches his arms beside it, getting ready for a new day.
Scene 2: kitchen
Motions:
pouring cereal, drinking
Narration:
Teddy is in the kitchen for breakfast. He pours cereal into a bowl in the kitchen, and drinks the soup, preparing his morning meal.
Scene 3: backyard
Motions:
watering plants, examining flower
Narration:
Teddy spends some time outside in the backyard. He waters the garden plants and kneels down to examine a flower, enjoying the fresh air and tending to the plants.
Scene 4: park
Motions:
jogging, breathing
Narration:
Teddy visits the park for some exercise. He jogs in the park, then he becomes very tired so he breaths heavily, regaining his energy.
Scene 5: forest
Motions:
hiking, sitting
Narration:
Teddy goes hiking on a forest trail. He hikes up the trail and then sits on a rock to rest, appreciating the tranquility of nature.
Scene 6: living room
Motions:
sitting, reclining
Narration:
Teddy winds down at the end of the day in the living room. He sits on the couch and then reclines to relax, ending his day in comfort and peace.
[Input] Mermaid’s Adventure
Listing 2: The example of the output from the high-level planning prompt, LABEL:lst:hlplan. We highlight scene, motions, narrations using same colors as Section 3.1.
[Output]
Scene 1: Coral Reef
Motions:
swimming, touching
Narration:
The Mermaid begins her day through the vibrant coral reef. She swims smoothly between colorful corals and touches the corals with her curiosity.
Scene 2: Sunken Ship
Motions:
diving, exploring
Narration:
The Mermaid diving toward the old, mysterious sunken ship, exploring its hidden corners. Her movements are careful as she examines every detail, wondering what secrets the ship might reveal.
Scene 3: Open Ocean
Motions:
gliding, waving
Narration:
In the open ocean, the Mermaid glides effortlessly through the water. She encounters a friendly dolphin and waves as it playfully circles around her before swimming off.
Scene 4: Kelp Forest
Motions:
wandering, inspecting
Narration:
The Mermaid enters a dense kelp forest, wandering through the tall, flowing plants. She pauses to inspect some unique shells caught on a kelp strand.
Scene 5: Sea Cave
Motions:
swimming, picking up
Narration:
The Mermaid ventures into a shadowy sea cave, swimming cautiously. She finds a glowing stone on the cave floor and picks it up, mesmerized by its mysterious light.
Scene 6: Lagoon
Motions:
lying, splashing
Narration:
The Mermaid finishes her journey by a secluded lagoon, lying near the shore. With a playful motion, she splashes water around, basking in the warm sunlight to end her day.
Listing 3: LLM prompt for fine-grained frame-level plan (Section 3.1, Scene-Level Fine-Grained Planning) per narration. We highlight Background, [entity, motion, caption] and layout [left, top, right, bottom] using same colors as Section 3.1.
Assuming the frame size is normalized to the range 0-1, you need to give a possible 6-frame layout plan at 1fps with the relevant regions, containing entity with corresponding motion, caption and bounding box involved in the input example motion and narration. You should follow these instructions:
1. [Background and Regions] You need to give a background of the videos, then list all regions with related entity, motion, caption, and bounding box, for each frame.
2. [Bounding Box Size] Each bounding box of the region is one rectangle or square box in the layout, and the size of boxes should be AS LARGE AS POSSIBLE. The width and height of each bounding box should be at least 0.2.
3. [Bounding Box Format] Every region should contain one related motion and caption, with the bounding boxes in the format of [[entity1, motion1, caption1], [left, top, right, bottom]]. If the entity doesnt involve any motion, use "none" as the motion (e.g. ["table", "none", "a table in the forest"]).
4. [Captions and Motions)] "IMPORTANT" For each entity, you should give a caption containing the entity and motion, with the provided background.
5. [Allowing Overlaps for Interaction Contexts] Regions can overlap if necessary, particularly when entities are interacting. For example, in "Teddy is pouring water into a bowl," you can have a region for "a bowl" overlapping or near Teddys region. Similarly, in "Teddy is sitting by the river," you can have a region for "river bank" overlapping or positioned beneath Teddys region.
5. [Allowing Overlaps for Interaction Contexts] Regions can overlap if necessary, particularly when entities are interacting. For example, in "Teddy is pouring water into a bowl," you can have a region for "a bowl" overlapping or near Teddys region. Similarly, in "Teddy is sitting by the river," you can have a region for "river bank" overlapping or positioned beneath Teddys region.
6. [Interaction with Objects] For an object listed in previous frames, when the entity is interacting with it in the current frames, you can still list the interacted object, then also use a whole region describing the character interacting with the object (e.g. [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]] before entity Teddy is interacting with it, [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]], [["Teddy", "sitting", "Teddy is sitting on the rock in the forest"], [0.0, 0.0, 0.4, 1.0]] when interacting).
7. [Static Interaction Objects] In most cases, the interacted objects should be static. For example, "Teddy is greeting to people" then "people" should have no bounding box changes. For some non-static objects like "the bottle" of "Teddy is drinking water", ignore such objects to be listed seperatedly.
8. [Motion Limitation] You should not use other motions outside the provided narration. The motion and caption should all be in PRESENT PROGRESSIVE.
9. [Motion Duration] One motion should last at least two frames.
10. [Reasoning] Add reasoning before you generate the region plan, explaining how you will allocate different events/motions to different frames and how the entity will be moving (e.g., left to right or staying static).
11. [Smart Motion Allocation] Allocate motions smartly by reasoning the frames they need (e.g., static motions like standing will require fewer frames but walking may need more). The background should not contain information about the characters (e.g., Teddys xxx is not allowed).
12. [Common Sense in Layout] Make sure the locations of the generated bounding boxes are consistent with common sense. You need to generate layouts from the close-up camera view of the event. The layout difference between two adjacent frames should not be too large, considering the small interval.
13. [Motions for Large Position Changes] If you want to move the entity largely by changing its bounding boxes (e.g., from most right to most left), make sure the motion naturally involves position changes (e.g., walking, running, flying, riding a bike). Otherwise, avoid big changes in related bounding boxes (e.g., cooking, jumping, playing guitar, etc., where bounding boxes should not change significantly).
14. [Big Regions Preference] We prioritize using large regions for main entities. For example, [0.2, 0.0, 0.8, 1.0] for a main teddy bear, as long as it fits the scene. Do not use small regions (like with only 0.2~0.3 width and height) as possible as you can.
15. [Motion Caption Consistency] Ensure that each caption accurately reflects the stated motion. For example, if the motion is "walking" the caption should match this exactly, such as "Teddy is walking in the park." Avoid any inconsistencies where the motion described in the caption does not match the stated motion, such as "Teddy is running in the park" when the motion is "walking." This consistency maintains clarity and alignment with the narration.
Use format:
*Reasoning*
reason
*Plan*
Background: background
Frame_1: [[entity1, motion1, caption1], [left, top, right, bottom]], [[entity2, motion2, caption2], [left, top, right, bottom]], ..., [[entity3, motion3, caption3], [left, top, right, bottom]]
Frame_2: [[entity1, motion1, caption1], [left, top, right, bottom]], [[entity2, motion2, caption2], [left, top, right, bottom]], ..., [[entity3, motion3, caption3], [left, top, right, bottom]]
Frame_6: [[entity1, motion1, caption1], [left, top, right, bottom]], [[entity2, motion2, caption2], [left, top, right, bottom]], ..., [[entity3, motion3, caption3], [left, top, right, bottom]]
Reasoning: ...
Example 1:
[Input]
Motion:
walking, sitting
Narration:
Teddy goes to a forest. He walks on the trail and then sits on a rock to rest, appreciating the tranquility of nature.
[Output]
*Reasoning*
Listed motions are: walking, sitting.
Related entities and motions: Teddy (from walking to sitting), rock (will be sit by Teddy in the end, but no related motions).
Motion frames allocation: The main entity involve motion changes is Teddy. Teddys motion changes from walking to sitting. As sitting takes less time, we should allocate more frames to walking. so the plan is 4 frames for walking and 2 frames for sitting. (Following bullet point 11 [Smart Motion Allocation])
Bounding box changes: For Teddy, the first motion is walking, which involves position changes a little bit. And the second motion is sitting, which should be a bounding-box-static motion. So the Teddy bear can have bounding boxes changes at first and finally interacts with the For the rock, it doesnt involve any motion, so it should be static from far from to close to, and finally interacting with Teddy. (Following bullet point 13 [Motions for Large Position Changes])
Interaction: This input narration involve interaction between the rock and Teddy and happens in late frames, so the rock and Teddy will have one merged-region where the entity is Teddy and the caption is about Teddy sitting on the rock. ALso the rock will also be listed from frame 1 to 3, while maintained from 4 to 6, as a static object. (Following bullet point 5 [Allowing Overlaps for Interaction Contexts], 6 [Interaction with Objects], 7 [Static Interaction Objects])
In conclusion, the plan can be: Teddy is at the right side at the beginning of the video, and theres also a rock on the left right corner. From Frame 1 to Frame 4, the Teddy bear is hiking on the trail moving from right to left to approaching the rock. In Frame 5 and Frame 6, it reaches the rock and sits on the rock (which is on the left corner) to rest, in the left side of the video.
*Plan*
Background: the forest
Frame_1: [["Teddy", "hiking", "Teddy is hiking on a trail in the forest"], [0.6, 0.0, 1.0, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
Frame_2: [["Teddy", "hiking", "Teddy is hiking on a trail in the forest"], [0.47, 0.0, 0.87, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
Frame_3: [["Teddy", "hiking", "Teddy is hiking on a trail in the forest"], [0.33, 0.0, 0.73, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
Frame_4: [["Teddy", "hiking", "Teddy is hiking on a trail in the forest"], [0.2, 0.0, 0.6, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
Frame_5: [["Teddy", "sitting", "Teddy is sitting on the rock in the forest"], [0.0, 0.0, 0.4, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
Frame_6: [["Teddy", "sitting", "Teddy is sitting on the rock in the forest"], [0.0, 0.0, 0.4, 1.0]], [["rock", "none", "a rock in the forest"], [0.0, 0.8, 0.2, 1.0]]
[Input]
Motions:
swimming, touching
Narration:
The Mermaid begins her day through the vibrant coral reef. She swims smoothly between colorful corals and touches the corals with her curiosity.
Listing 4: The example output of fine-grained planning prompt LABEL:lst:fgplan. We highlight Background, [entity, motion, caption] and layout [left, top, right, bottom] using same colors as Section 3.1.
[Output]
*Reasoning*
Listed motions are: swimming, touching.
Related entities and motions: Mermaid (from swimming to touching), corals (being touched by the mermaid).
Motion frames allocation: The main entity involving motion changes is the Mermaid. The first motion, swimming, can last relatively longer with more frame coverage as it involves position changes. Touching corals happens afterward and should occupy fewer frames with minimal position changes.
Bounding box changes: The swimming motion involves position changes, so the bounding box for the Mermaid can have slight shifts, swimming through the corals from left to right. The touching motion of the Mermaid should be static, focusing on the interaction at the coral, which is near the center of the frame.
Interaction: The interaction between the Mermaid and the corals happens later in the narration, so a combined region featuring the Mermaid touching the corals will be involved in the last two frames.
Non-interacting corals will appear in all frames, having stable bounding boxes across the video to show the environment and the Mermaids interaction context.
The plan is as follows: The Mermaid swims through the vibrant coral reef, starting from near the left, moving towards the center, where the corals are located. On reaching the center, she begins touching the corals with fascination.
*Plan*
Background: the vibrant coral reef full of colors and life
Frame_1: [["Mermaid", "swimming", "The Mermaid is swimming through the vibrant coral reef"], [0.0, 0.0, 0.4, 1.0]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]
Frame_2: [["Mermaid", "swimming", "The Mermaid is swimming through the vibrant coral reef"], [0.15, 0.0, 0.55, 1.0]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]
Frame_3: [["Mermaid", "swimming", "The Mermaid is swimming through the vibrant coral reef"], [0.3, 0.0, 0.7, 1.0]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]
Frame_4: [["Mermaid", "swimming", "The Mermaid is swimming through the vibrant coral reef"], [0.45, 0.0, 0.85, 1.0]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]
Frame_5: [["Mermaid", "touching", "The Mermaid is touching the corals with her curiosity"], [0.4, 0.1, 0.7, 0.9]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]
Frame_6: [["Mermaid", "touching", "The Mermaid is touching the corals with her curiosity"], [0.4, 0.1, 0.7, 0.9]], [["corals", "none", "Colorful corals in the reef"], [0.5, 0.3, 0.8, 0.6]]